Solve 7x – 12 = 16Solve 7x – 12 = 16Solve 7x – 12 = 16

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

7x -12 = 16

7x = 16 + 12

x = 28/7

x = 4


Thanks